In a reversible adiabatic manner, 17.6 m3 /min of air are compressed from 277K and 101 kPa to 700 kPa. Determine the power required.

W =


1
Expert's answer
2021-10-08T15:44:04-0400

"\frac{p_1}{p_2}=(\frac{V_2}{V_1})^k,"

"W=\frac{p_1V_1}{60(k-1)}(1-(\frac{p_2}{p_1})^{\frac{k-1}k})=3.97~\frac{kW}s."